# Amortized Variational Transdimensional Inference

### v1

@@ D.3 Monte Carlo estimation of information
Appendix E Robust variable selection example details and additional results
The likelihood setup is
$p\bigl(y_{i}\mid\bm{x}_{i},\bm{\beta},\bm{\gamma}\bigr)$ $=\bm{\mu}(\bm{x}_{i})+(1-\alpha)\,\mathcal{N}\!\Bigl(y_{i};0,\sigma_{1}^{2}\Bigr)\;+\;\alpha\,\mathcal{N}\!\Bigl(y_{i};0,\sigma_{2}^{2}\Bigr),$ (58)
@@ Appendix E Robust variable selection example details and additional results
E.1 Sweep of increasing cardinality
Using the focused prior setup on both the medium and high misspecification level targets, we sweep the cardinality of the model space $|\pazocal{M}|$ from $2^{9}=512$ to $2^{24}=16,777,216$ and compute the cross entropy $H(\pi,q_{\psi,\phi})$. The reversible jump MCMC process in Appendix E.4 was used to obtain gold 1,000 standard samples from the “true” $\pi(m,\bm{\theta}_{m})$, based on 50,000 chain iterations and retaining every 50th sample. The results of 1,024 replicate chains are visualized in Figure 4, which shows how the cross entropy $H(\pi,q_{\psi,\phi})$ changes with $|\pazocal{M}|$. Although $H(\pi,q_{\psi,\phi})$ generally increases with $|\pazocal{M}|$, we notice that, as expected (with the existing tuning parameters), the surrogate method (blue bars) performs comparably with the other methods for the smaller model spaces ($|\pazocal{M}|=2^{9}$), whereas the neural density (orange bars) performs consistently as $|\pazocal{M}|$ increases.
Figure 4: A simulation study showing the cross entropy (NLL) between reversible jump MCMC samples and an amortized variational transdimensional density using rational quadratic spline CoSMIC flows under a fixed number of iterations (30,000). Each cardinality was run with 10 independently sampled synthetic data sets.
E.2 Focused versus wide priors
@@
For the $\sigma_{\beta}=10$ wide prior setting (Figures 6, 8, 10) performance at first glance appears much worse, particularly in terms of estimating model probabilities. However, on closer inspection this is not the case. It is well known (e.g. Gelman and Yao, (2020)) that the marginal likelihood (a.k.a. model evidence; a component of the posterior model probability) can be highly sensitive to diffuse priors. In such cases (as with $\sigma_{\beta}=10$) the posterior will tend to unreasonably favour those models with fewer parameters, and particularly (in the case of regression models) the null model with no predictors, even in the presence of a very clear relationship between predictors and response. This effect can be clearly seen in Figures 6, 8, 10 (top row), where the null model (indicated by a circle) is given far higher posterior model probability on the $\pi(m)$ axis than the actual data generating process (triangles). In contrast, the true data generating process (triangles) is generally given a high posterior model probability (comparable with the focused prior setting in Figures 5, 7, 9) under the VTI approximation. From these results we conclude that: (i) the posterior model probabilities that depend on the marginal likelihood (i.e., the estimates of $\pi(m)$ on the $x$-axis) have been affected by the wide prior to unreasonably favour models with less parameters; (ii) the VTI-based posterior model probability estimates suggest that they are less sensitive to the undesirable effects of this prior; and (iii) in combination the resulting plots in Figures 6, 8, 10 (top row) only appear to indicate worse performance of VTI compared to the gold standard than is actually the case.
E.3 Within model comparison
@@
Figure 11: Bivariate plot comparison between reversible jump MCMC (left) and VTI (right) using spline flow composition of four layers and six blocks on the first synthetic high-misspecification data set from the Figure 2 (main text) example.
E.4 Reversible jump MCMC for robust variable selection
@@
Birth Move $(m,\beta)\rightarrow(m^{\prime},\beta)$: A birth move flips a bit from 0 to 1. The acceptance ratio is derived as follows:
$\alpha_{\text{birth}}=\min\left\{1,\frac{p(y\mid\beta,m^{\prime})\,p(\beta\mid m^{\prime})\,\pi(m^{\prime})}{p(y\mid\beta,m)\,p(\beta\mid m)\,\pi(m)}\right\}$
Death Move $(m,\beta)\rightarrow(m^{\prime},\beta)$:
A death move flips a bit from 1 to 0. The acceptance ratio is the reciprocal of the birth ratio $\alpha_{\text{death}}=\frac{1}{\alpha_{\text{birth}}}$.
@@
Thus, the Jacobian does not affect the acceptance ratio.
@@
$\alpha_{\text{within}}=\min\left\{1,\frac{p(y\mid\beta^{\prime},m)\,p(\beta^{\prime}\mid m)}{p(y\mid\beta,m)\,p(\beta\mid m)}\right\}$
Since the proposal is symmetric, the proposal densities cancel out in the acceptance ratio.
Appendix F Example description: Bayesian inference of multi-layer-perceptron directed acyclic graph discovery
@@
$\bm{X}$ $\in\mathbb{R}^{N_{d}\times n}$ (columns are i.i.d. samples)
$\bm{P}$ $\in\mathcal{P}_{N_{d}}$ permutation matrix (node order)
$\bm{U}$ $\in\{0,1\}^{N_{d}\times N_{d}}$ strictly upper–triangular edge mask
$\bm{A}$ $=\bm{P}^{\top}\bm{U}\bm{P}$ adjacency in canonical order (code default)
$\operatorname{pa}_{\bm{A}}\!\left(j\right)$ $=\{\,i<j:\bm{U}_{ij}=1\,\}$ parents of node $j$ in the sorted order.
@@
Fix hidden width $H$ and a model indicator $m=(\bm{P},\bm{U})$. For each non-root node $j=2,\dots,N_{d}$ define parameters
$\bm{\theta}^{(j)}=\bigl(W^{(1)}_{j},b^{(1)}_{j},W^{(2)}_{j},b^{(2)}_{j}\bigr)\in\mathbb{R}^{(j+2)H+1},$
with $W^{(1)}_{j}\!\in\!\mathbb{R}^{H\times(j-1)},\;b^{(1)}_{j}\!\in\!\mathbb{R}^{H},\;W^{(2)}_{j}\!\in\!\mathbb{R}^{1\times H},\;b^{(2)}_{j}\!\in\!\mathbb{R}$. Let $\bm{u}_{j}:=\bm{U}_{1:(j-1),\,j}$ be the (j1)(j-1)-vector of active parents. For $\bm{x}=(x_{1},\dots,x_{j-1})^{\!\top}$,
$f_{j}(\bm{x};\bm{\theta}^{(j)},\bm{U})=W^{(2)}_{j}\,\operatorname{ReLU}\!\bigl(W^{(1)}_{j}(\bm{x}\odot\bm{u}_{j})+b^{(1)}_{j}\bigr)+b^{(2)}_{j},\qquad f_{1}(\,\cdot\,)\equiv 0.$
@@
Let $\varpi$ be the permutation associated with $\bm{P}$ (so $x_{\varpi(j)}$ is column $j$ after sorting). With homoscedastic noise $\sigma^{2}$,
$\boxed{\log p\!\bigl(\bm{X}\!\mid\!\bm{P},\bm{U},\bm{\theta}\bigr)=-\frac{n\!N_{d}}{2}\log\!\bigl(2\pi\sigma^{2}\bigr)-\frac{1}{2\sigma^{2}}\sum_{s=1}^{n}\sum_{j=1}^{N_{d}}\Bigl(x_{\varpi(j)}^{(s)}-f_{j}\!\bigl(\bm{x}_{\varpi(1:j-1)}^{(s)};\bm{\theta}^{(j)},\bm{U}\bigr)\Bigr)^{2}}$
@@
$\boxed{p(\bm{\theta}\!\mid\!\bm{P},\bm{U})=\!\!\prod_{k\in C(m)}\mathcal{N}\!\bigl(\bm{\theta}_{k};0,\sigma_{0}^{2}\bigr)}$
@@
$\boxed{p(\bm{P},\bm{U})\;\propto\;\exp\!\bigl(-\lambda\,\|\bm{U}\|_{1}\bigr)},\qquad\lambda\geq 0,$
with $\bm{P}$ a permutation matrix and $\bm{U}$ strictly upper triangular.
@@
F.1 Data generating process
The data generating procedure generally follows the simulation design in Thompson et al., (2025).
@@ F.1 Data generating process
$\bm{P}$ $\sim\text{Uniform}\bigl(\mathcal{P}_{N_{d}}\bigr),$
$\bm{U}_{ij}$ $\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\operatorname{Bernoulli}(\rho_{\operatorname{Edge}}),\hskip 18.49988pt1\leq i<j\leq N_{d},$
$\bm{A}$ $=\bm{P}^{\top}\bm{U}\bm{P}.$
@@
$\bm{\theta}^{(j)}=\bigl(W^{(1)}_{j},\;\beta\,b^{(1)}_{j},\;W^{(2)}_{j},\;\beta\,b^{(2)}_{j}\bigr),\qquad\bigl[\bm{\theta}^{(j)}\bigr]_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}[-0.7,-0.3]\cup[0.3,0.7],$
@@
For $m=(\bm{P},\bm{U})$, $C(m)=C(\bm{U})\subseteq\{1,\ldots,\dim\bm{\theta}\}$ keeps exactly the coordinates satisfying the conditions:
@@
    Column $i$ of $W^{(1)}_{j}$ is active iff $\bm{U}_{ij}=1$;
@@
    If $\sum_{i<j}\bm{U}_{ij}=0$ then all parameters in $\bm{\theta}^{(j)}$ are masked.
(The permutation $\bm{P}$ has no effect on the mask.)
@@
Let $\varpi$ be the permutation induced by $\bm{P}$. For each sample $s=1,\dots,n$ generate sequentially
$x_{\varpi(1)}^{(s)}$ $=\sigma\,\varepsilon_{1s},$
$x_{\varpi(j)}^{(s)}$ $=f_{j}\!\bigl(\bm{x}_{\varpi(1:j-1)}^{(s)};\bm{\theta}^{(j)},\bm{U}\bigr)+\sigma\,\varepsilon_{js},\qquad j=2,\dots,N_{d},$
@@
$f_{j}(\bm{z};\bm{\theta}^{(j)},\bm{U})=W^{(2)}_{j}\,\operatorname{ReLU}\!\bigl(W^{(1)}_{j}(\bm{z}\odot\bm{u}_{j})+\beta\,b^{(1)}_{j}\bigr)+\beta\,b^{(2)}_{j},\quad\bm{u}_{j}:=\bm{U}_{1:(j-1),\,j}.$
@@
$\bm{X}=\begin{bmatrix}\bm{x}^{(1)}\\[-2.0pt] \vdots\\[-2.0pt] \bm{x}^{(n)}\end{bmatrix}\in\mathbb{R}^{n\times N_{d}},\quad\text{stored in topological order }\bigl(x_{\varpi(1)},\dots,x_{\varpi(N_{d})}\bigr).$
F.2 Comparison metrics
Given knowledge of a “true” adjacency matrix $A$, each experiment uses four scores for comparison with the estimated posterior: F1, structured Hamming distance (SHD), Brier score, and area under the receiver operating characteristic curve (AUROC). This follows the experiment setup in Thompson et al., (2025).
F.3 Common inference setup
For each data set in both the simulation study and real data example, VTI is run a total of 10 replicates using different random seeds, and the posterior is selected where the terminal loss is minimized. For DAGMA, the sparsity hyperparameter is swept from $\lambda^{\operatorname{min}}=10^{-3}$ to $\lambda^{\operatorname{max}}=1$ over 10 logarithmically spaced values. For the autoregressive flow, we use Affine(5,5) (see Section˜A.2) with a context encoder designed as follows:
$\delta(\bm{P},\bm{U})=\sigma^{\lceil\times 2\rceil}\circ\dots\circ\sigma^{\lceil\times 2\rceil}\circ(\bm{P}^{\top}\bm{U}\bm{P}),$
where $\sigma^{\lceil\times 2\rceil}(x):=Wx+b$ broadcasts from $|x|$ to the first power of 2 greater than or equal to $2|x|$. The final dimension of $\delta(\bm{P},\bm{U})$ is 4096.
F.4 Simulation design

### v2

@@ E.3 Monte Carlo estimation of information
Appendix F Robust variable selection example details and additional results
The likelihood is
$p\bigl(y_{i}\mid\bm{x}_{i},\bm{\beta},\bm{\gamma}\bigr)$ $=(1-\alpha)\,\mathcal{N}\!\Bigl(y_{i};\bm{\mu}(\bm{x}_{i}),\sigma_{1}^{2}\Bigr)\;+\;\alpha\,\mathcal{N}\!\Bigl(y_{i};\bm{\mu}(\bm{x}_{i}),\sigma_{2}^{2}\Bigr),$ (54)
@@ Appendix F Robust variable selection example details and additional results
F.1 Focused versus wide priors
@@
For the $\sigma_{\beta}=10$ wide prior setting (Figures 6, 8, 10) performance at first glance appears much worse, particularly in terms of estimating model probabilities. However, on closer inspection this is not the case. It is well known (e.g. Gelman and Yao [2020]) that the marginal likelihood (a.k.a. model evidence; a component of the posterior model probability) can be highly sensitive to diffuse priors. In such cases (as with $\sigma_{\beta}=10$) the posterior will tend to unreasonably favour those models with fewer parameters, and particularly (in the case of regression models) the null model with no predictors, even in the presence of a very clear relationship between predictors and response. This effect can be clearly seen in Figures 6, 8, 10 (top row), where the null model (indicated by a circle) is given far higher posterior model probability on the $\pi(m)$ axis than the actual data generating process (triangles). In contrast, the true data generating process (triangles) is generally given a high posterior model probability (comparable with the focused prior setting in Figures 5, 7, 9) under the VTI approximation. From these results we conclude that: (i) the posterior model probabilities that depend on the marginal likelihood (i.e., the estimates of $\pi(m)$ on the $x$-axis) have been affected by the wide prior to unreasonably favour models with less parameters; (ii) the VTI-based posterior model probability estimates suggest that they are less sensitive to the undesirable effects of this prior; and (iii) in combination the resulting plots in Figures 6, 8, 10 (top row) only appear to indicate worse performance of VTI compared to the gold standard than is actually the case.
F.2 Within model comparison
@@
Figure 11: Multivariate plot comparison between reversible jump MCMC (left) and VTI (right) using spline flow composition of four layers and six blocks on the first synthetic narrow-prior high-misspecification data set from the Figure 2 (main text) example.
F.3 Baseline reversible jump MCMC for robust variable selection
@@
Thus, the Jacobian does not affect the acceptance probability.
Birth Move $(m,\beta)\rightarrow(m^{\prime},\beta)$: A birth move flips a bit in $m\rightarrow m^{\prime}$ from 0 to 1. Given the birth/death move ratio
$r_{\text{b-d}}(m,m^{\prime},\beta)=\frac{p(y\mid\beta,m^{\prime})\,p(\beta\mid m^{\prime})\,\pi(m^{\prime})}{p(y\mid\beta,m)\,p(\beta\mid m)\,\pi(m)},$
the acceptance probability is:
$\alpha_{\text{birth}}(m,m^{\prime},\beta)=\min\left\{1,r_{\text{b-d}}(m,m^{\prime},\beta)\right\}.$
Death Move $(m,\beta)\rightarrow(m^{\prime},\beta)$: A death move flips a bit in $m\rightarrow m^{\prime}$ from 1 to 0. Using the same birth/death move ratio, the acceptance probability is:
$\alpha_{\text{death}}(m,m^{\prime},\beta)=\min\left\{1,r_{\text{b-d}}(m,m^{\prime},\beta)\right\}.$
@@
$\alpha_{\text{within}}(m,\beta,\beta^{\prime})=\min\left\{1,\frac{p(y\mid\beta^{\prime},m)\,p(\beta^{\prime}\mid m)}{p(y\mid\beta,m)\,p(\beta\mid m)}\right\}$
Since the proposal is symmetric, the proposal densities cancel out in the acceptance probability.
Appendix G Example description: Bayesian inference of multi-layer-perceptron directed acyclic graph discovery
@@
$N_{d}$ number of nodes in graph
$n$ number of data samples
$\bm{X}$ $\in\mathbb{R}^{n\times N_{d}}$ (rows are i.i.d. samples)
$\mathbf{P}$ $\in\mathcal{P}_{N_{d}}$ permutation matrix (node order)
$\mathbf{U}$ $\in\{0,1\}^{N_{d}\times N_{d}}$ strictly upper–triangular edge mask
$\mathbf{A}$ $=\mathbf{P}^{\top}\mathbf{U}\mathbf{P}$ adjacency in canonical order (code default)
$\operatorname{pa}_{\mathbf{A}}\!\left(j\right)$ $=\{\,i<j:\mathbf{U}_{ij}=1\,\}$ parents of node $j$ in the sorted order.
@@
Fix hidden width $H$ and a model indicator $m=(\mathbf{P},\mathbf{U})$. For each non-root node $j=2,\dots,N_{d}$ define parameters
$\bm{\theta}^{(j)}=\bigl(\mathbf{W}^{(1)}_{j},b^{(1)}_{j},\mathbf{W}^{(2)}_{j},b^{(2)}_{j}\bigr)\in\mathbb{R}^{(j+2)H+1},$
with $\mathbf{W}^{(1)}_{j}\!\in\!\mathbb{R}^{H\times(j-1)},\;b^{(1)}_{j}\!\in\!\mathbb{R}^{H},\;\mathbf{W}^{(2)}_{j}\!\in\!\mathbb{R}^{1\times H},\;b^{(2)}_{j}\!\in\!\mathbb{R}$. Let $\bm{u}_{j}:=\mathbf{U}_{1:(j-1),\,j}$ be the (j1)(j-1)-vector of active parents. Writing $\bm{X}_{1:j-1}$ to denote the $1,\dots,j-1$ columns of $\bm{X}$,
$f_{j}(\bm{X}_{1:j-1};\bm{\theta}^{(j)},\mathbf{U})=\mathbf{W}^{(2)}_{j}\,\operatorname{ReLU}\!\bigl(\mathbf{W}^{(1)}_{j}(\bm{X}_{1:j-1}\odot\bm{u}_{j})+b^{(1)}_{j}\bigr)+b^{(2)}_{j},\hskip 18.49988ptf_{1}(\,\cdot\,)\equiv 0.$ (55)
@@
Let $\varpi$ be the permutation associated with $\mathbf{P}$ (so $\bm{X}_{\varpi(j)}$ is column $j$ after sorting). With homoscedastic noise $\sigma^{2}$,
$\boxed{\log p\!\bigl(\bm{X}\!\mid\!\mathbf{P},\mathbf{U},\bm{\theta}\bigr)=-\frac{n\!N_{d}}{2}\log\!\bigl(2\pi\sigma^{2}\bigr)-\frac{1}{2\sigma^{2}}\sum_{s=1}^{n}\sum_{j=1}^{N_{d}}\Bigl(\bm{X}_{\varpi(j)}^{(s)}-f_{j}\!\bigl(\bm{X}_{\varpi(1:j-1)}^{(s)};\bm{\theta}^{(j)},\mathbf{U}\bigr)\Bigr)^{2}}$
@@
$\boxed{p(\bm{\theta}\!\mid\!\mathbf{P},\mathbf{U})=\!\!\prod_{k\in C(m)}\mathcal{N}\!\bigl(\bm{\theta}_{k};0,\sigma_{0}^{2}\bigr)}$
@@
$\boxed{p(\mathbf{P},\mathbf{U})\;\propto\;\exp\!\bigl(-\lambda\,\|\mathbf{U}\|_{1}\bigr)},\qquad\lambda\geq 0,$
with $\mathbf{P}$ a permutation matrix and $\mathbf{U}$ strictly upper triangular.
@@
G.1 Data generating process
The data generating procedure generally follows the simulation design in Thompson et al. [2025].
@@ G.1 Data generating process
$\mathbf{P}$ $\sim\text{Uniform}\bigl(\mathcal{P}_{N_{d}}\bigr),$
$\mathbf{U}_{ij}$ $\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\operatorname{Bernoulli}(\rho_{\operatorname{Edge}}),\hskip 18.49988pt1\leq i<j\leq N_{d},$
$\mathbf{A}$ $=\mathbf{P}^{\top}\mathbf{U}\mathbf{P}.$
@@
$\bm{\theta}^{(j)}=\bigl(\mathbf{W}^{(1)}_{j},\;\beta\,b^{(1)}_{j},\;\mathbf{W}^{(2)}_{j},\;\beta\,b^{(2)}_{j}\bigr),\qquad\bigl[\bm{\theta}^{(j)}\bigr]_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}[-0.7,-0.3]\cup[0.3,0.7],$
@@
For $m=(\mathbf{P},\mathbf{U})$, $C(m)=C(\mathbf{U})\subseteq\{1,\ldots,\dim\bm{\theta}\}$ keeps exactly the coordinates satisfying the conditions:
@@
    Column $i$ of $\mathbf{W}^{(1)}_{j}$ is active iff $\mathbf{U}_{ij}=1$;
@@
    If $\sum_{i<j}\mathbf{U}_{ij}=0$ then all parameters in $\bm{\theta}^{(j)}$ are masked.
(The permutation $\mathbf{P}$ has no effect on the mask.)
@@
Let $\varpi$ be the permutation induced by $\mathbf{P}$. For each sample $s=1,\dots,n$ generate sequentially
$\bm{X}_{\varpi(1)}^{(s)}$ $=\sigma\,\varepsilon_{1s},$
$\bm{X}_{\varpi(j)}^{(s)}$ $=f_{j}\!\bigl(\bm{X}_{\varpi(1:j-1)}^{(s)};\bm{\theta}^{(j)},\mathbf{U}\bigr)+\sigma\,\varepsilon_{js},\qquad j=2,\dots,N_{d},$
@@
$f_{j}(\bm{z};\bm{\theta}^{(j)},\mathbf{U})=\mathbf{W}^{(2)}_{j}\,\operatorname{ReLU}\!\bigl(\mathbf{W}^{(1)}_{j}(\bm{z}\odot\bm{u}_{j})+\beta\,b^{(1)}_{j}\bigr)+\beta\,b^{(2)}_{j},\quad\bm{u}_{j}:=\mathbf{U}_{1:(j-1),\,j}.$
@@
$\bm{X}=\begin{bmatrix}\bm{X}^{(1)}\\[-2.0pt] \vdots\\[-2.0pt] \bm{X}^{(n)}\end{bmatrix}\in\mathbb{R}^{n\times N_{d}},\quad\text{stored in topological order }\bigl(\bm{X}_{\varpi(1)},\dots,\bm{X}_{\varpi(N_{d})}\bigr).$
G.2 Comparison metrics
Given knowledge of a “true” adjacency matrix $A$, each experiment uses four scores for comparison with the estimated posterior: F1, structured Hamming distance (SHD), Brier score, and area under the receiver operating characteristic curve (AUROC). This follows the experiment setup in Thompson et al. [2025].
G.3 Common inference setup
For each data set in both the simulation study and real data example, VTI is run a total of 10 replicates using different random seeds, and the posterior is selected where the terminal loss is minimized. For DAGMA, the sparsity hyperparameter is swept from $\lambda^{\operatorname{min}}=10^{-3}$ to $\lambda^{\operatorname{max}}=1$ over 10 logarithmically spaced values. For the autoregressive flow, we use Affine(5,5) (see Appendix A.2) with a context encoder designed as follows:
$\delta(\mathbf{P},\mathbf{U})=\sigma^{\lceil\times 2\rceil}\circ\dots\circ\sigma^{\lceil\times 2\rceil}\circ(\mathbf{P}^{\top}\mathbf{U}\mathbf{P}),$
where $\sigma^{\lceil\times 2\rceil}(x):=Wx+b$ broadcasts from $|x|$ to the first power of 2 greater than or equal to $2|x|$. The final dimension of $\delta(\mathbf{P},\mathbf{U})$ is 4096.
G.4 Simulation design